# Arithmetical Complexity and Absoluteness of Rigidity Phenomena for Ulam Sequences

### v1

@@ 8 Further directions and open problems
The framework developed here applies to any additive construction whose membership relation is recursive and whose conjectured large-scale structure can be encoded by finite patterns with linear dependence on parameters. Beyond classical Ulam sequences and higher-dimensional Ulam sets, there are a number of natural candidates.
@@
  • Constructions in non-standard ambient groups, such as ordered abelian groups other than $(\mathbb{Z},+)$, or in semigroups with additional combinatorial constraints (e.g. restricted sumsets). The arithmetisation methods of Sections 23 could extend with only notational changes to such settings, as long as the underlying group and ordering are recursively presented. It would be interesting to carry out a similar complexity and absoluteness analysis for these multidimensional Ulam sets, and to compare the resulting expansion theories with known model-theoretic work on expansions of $(\mathbb{Z}^{d},+)$.
@@
  • Polynomial Ulam sequences in the sense of Sheydvasser [Sheydvasser2021LinearPolynomials], and more general algebraic deformations where the parameters $(a,b)$ are replaced by integer polynomials or higher-dimensional algebraic data. For example, one can ask whether Ulam-type constructions in ordered rings such as $\mathbb{Z}[X]$ or $\mathbb{Z}[X_{1},\dots,X_{r}]$ admit similar finite pattern codings and corresponding arithmetical complexity bounds.

### v2

@@ 8 Further directions and open problems
The framework developed here applies to any additive construction whose membership relation is recursive and whose conjectured large-scale structure can be encoded by finite patterns with linear dependence on parameters. Natural candidates for similar analysis include:
  • Higher-dimensional Ulam sets in $\mathbb{Z}^{n}$, as studied by Kravitz–Steinerberger [KravitzSteinerberger2018] and Hinman–Kuca–Schlesinger–Sheydvasser [HinmanKucaSchlesingerSheydvasser2019Involve]. The arithmetisation methods extend naturally to recursively presented ordered abelian groups.
@@
  • Polynomial Ulam sequences in the sense of Sheydvasser [Sheydvasser2021LinearPolynomials], and more general algebraic deformations where the parameters $(a,b)$ are replaced by integer polynomials or higher-dimensional algebraic data.